# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ Abstract
Index Terms:
Non-Terrestrial Networks (NTN), Space-Air-Ground Integrated Networks (SAGIN), Artificial Intelligence (AI), Machine Learning (ML), Deep Learning (DL), 5G-Advanced, 6G, Satellite, Beam-hopping, Handover, Spectrum sharing, Doppler shift, Resource allocation, Computational offloading, Network routing, Network slicing, Channel estimation, Security, Open Radio Access Network (O-RAN), RAN Intelligent Controller (RIC).I Introduction
@@
The Third Generation Partnership Project (3GPP) has already started the standardization towards the 5th Generation (5G)-Advanced in Release 17 and 18 to facilitate its worldwide deployment[1, 2]. 5G-Advanced provides much higher data rates, lower latency, increased capacity, and more efficient spectrum utilization than any of its predecessors. It supports a wide range of applications encompassing all 5G use cases such as Ultra Reliable Low Latency Communications (URLLC), massive Machine Type Communication (mMTC), and enhanced Mobile Broadband (eMBB) communication with different Key Performance Indicator (KPI) requirements [3]. Nevertheless, future applications such as Augmented Reality (AR), Virtual Reality (VR), Tactile Internet, Holographic Type Communication (HTC), remote health and surgery, etc. require extremely high throughput, low latency, high reliability, and ubiquity at the same time which cannot be met with current technological standards [4]. Consequently, the next-generation global wireless standard, namely, 6th Generation (6G) has become the current research focus for the industry and research community [5].
@@
6G is expected to provide an extremely high data rate (peak data rate up to 1 Tbps and user-experienced data rate up to 10 Gbps, around 100 times higher than 5G), very low latency (in the order of s), high reliability (around 100 times better than 5G) and extreme coverage to support the diverse set of future applications [6, 7, 8]. Due to the limited coverage area and geographical constraints, it is not possible to guarantee ubiquitous connectivity with existing terrestrial-only network infrastructures. Non-Terrestrial Networks (NTNs), networks involving space and aerial platforms, can provide us with multicast opportunities over very large areas as well as can serve users even in remote areas or during times of natural calamities [4]. Furthermore, the launching and maintenance costs for satellites have significantly decreased as they are deployed at lower heights (typically around 600 km). These satellites can provide much higher throughput and lower latency compared to legacy satellites and potentially can support different use cases of 6G. So NTN is considered to become one of the major technological enablers of future 6G networks visioning connectivity anywhere and anytime [4, 9, 10, 11]. Tech giants such as SpaceX Starlink, Amazon Kuiper, and OneWeb have already begun to invest billions of dollars in this field, reflecting its massive potential for future growth [12].
@@
Although NTN presents numerous potential benefits for the development of future 6G networks, it also entails several challenges that need to be addressed, primarily due to the unique characteristics of its mobility and propagation environments [13, 14]. Due to the long distances between the space-borne Base Stations (BS) and the ground User Equipment (UE), the propagation delay is usually higher in NTN environments. Additionally, high-speed air or space-borne platforms necessitate modifications to existing handover and paging protocols, as well as introduce a significant Doppler shift in carrier frequencies. The large path loss also increases the minimum power requirement for reliable transmission, initiating the need for novel beam and resource allocation strategies. Spectrum sharing in the same frequency band with existing terrestrial or other services also requires further study in order to avoid interference between terrestrial and non-terrestrial users. Even though currently there are some stand-alone satellite network deployments, the ultimate goal is the convergence of terrestrial and non-terrestrial environments for extreme network performance in 6G. [4]. This potential integrated environment requires efficient computing, routing, and slicing algorithms for meeting the expected KPI requirements of 6G.
@@
Artificial intelligence (AI) is currently having a profound and revolutionary impact on a multitude of industries, including but not limited to healthcare, military, transportation, and e-Commerce [15]. AI encompasses a wide array of smart machines, while Machine Learning (ML) is a popular subset of AI that allows machines to learn from large amounts of data and make decisions without the need for explicit programming [16]. Deep Learning (DL) is a special subset of ML that studies Artificial Neural Networks (ANNs) which contain more than one hidden layer, often implemented to simulate the human brain [17]. DL is currently being leveraged in various applications, such as computer vision, speech recognition, and bioinformatics, outperforming human-level performance in these particular domains. The cellular domain is still in its infancy in terms of AI integration [18] compared to other fields due to the complex and dynamic nature of wireless networks. As an integral part of 6G networks, challenges associated with NTN deployment provides an enticing field for AI applications. However, while deploying algorithms in a real environment, practical implementation difficulties may arise to provide reliable vertical connectivity between the ground and space networks. To reach optimal performance, theoretical advancements in communication system design must be complemented by appropriate AI solutions for NTN integration into 6G.
@@
3GPP |
Third Generation Partnership Project |
5G |
5th Generation |
6G |
6th Generation |
AC |
Actor-Critic |
ACO |
Ant-Colony based Optimization |
AI |
Artificial Intelligence |
ANN |
Artificial Neural Networks |
AoA |
Angle of Arrival |
AoD |
Angle of Departure |
AR |
Augmented Reality |
ARIMA |
Auto-Regressive Moving Average |
ARMA |
Auto-Regressive Integrated Moving Average |
ATM |
Asynchronous Transfer Mode |
BS |
Base Station |
CCI |
Co-Channel Interference |
CDMA |
Code Division Multiple Access |
CG |
Coordinate Graph |
CNN |
Convolutional Neural Network |
CSD |
Cyclo-Stationary Detection |
CSI |
Channel State Information |
CU |
Central Unit |
DBN |
Deep Belief Network |
DcL |
Decentralized Learning |
DDPG |
Deep Deterministic Policy Gradient |
DDQN |
Double Deep Q-Learning Network |
DL |
Deep Learning |
DN |
Deconvolutional Network |
DoS |
Denial-of-Service |
DP |
Dynamic Programming |
DPG |
Deterministic Policy Gradient |
DQN |
Deep Q-Learning Network |
DRL |
Deep Reinforcement Learning |
DRN |
Deep Residual Network |
DRQN |
Deep Recurrent Q-Learning Network |
DU |
Distributed Unit |
ED |
Energy Detection |
eMBB |
enhanced Mobile Broadband |
ELM |
Extreme Learning Machine |
ESA |
European Space Agency |
ESN |
Echo-State Network |
EVD |
Eigen Value-based Detection |
FCNN |
Fully Connected Neural Network |
FDMA |
Frequency Division Multiple Access |
FL |
Federated Learning |
FlexRIC |
Flexible RIC |
GA |
Genetic Algorithm |
GDM |
Generative Diffusion Model |
GMM |
Gaussian Mixture Model |
GAN |
Generative Adversarial Network |
GNN |
Graph Neural Network |
GEO/GSO |
Geostationary Earth Orbit |
GNSS |
Global Navigation Satellite System |
GPS |
Global Positioning System |
GRU |
Gated Recurrent Unit |
HAPS |
High Altitude Platform System |
HARQ |
Hybrid Automatic Repeat Request |
HIBS |
High-altitude International Mobile Base Station |
HTC |
Holographic Type Communication |
IC |
Integrated Circuit |
ICI |
Inter-Carrier Interference |
IoT |
Internet of Things |
IP |
Internet Protocol |
ISL |
Inter-Satellite Link |
ITU |
International Telecommunication Union |
KKT |
KarushKuhnTucker |
KNN |
k-Nearest Neighbor |
KPI |
Key Performance Indicator |
LEO |
Low Earth Orbit |
LoS |
Line of Sight |
LSM |
Liquid State Machine |
LSTM |
Long-Term Short Memory |
@@
MADRL |
Multi-Agent Deep Reinforcement Learning |
MAP |
Maximum A Posteriori |
MARL |
Multi-Agent Reinforcement Learning |
MC |
Monte Carlo method |
MDP |
Markov Decision Process |
MEMS |
Micro-Electro-Mechanical Systems |
MEO |
Medium Earth Orbit |
MiM |
Man-in-the-Middle |
ML |
Machine Learning |
MLP |
Multi-Layer Perceptron |
MMSE |
Minimum Mean Square Error |
mMTC |
massive Machine Type Communication |
MNL |
Minimum Network Load |
MSQ |
Maximum Signal Quality |
MST |
Maximum Service Time |
mULC |
massive Ultra-reliable low-Latency Communication |
NGEO/NGSO |
Non-Geostationary Earth Orbit |
NN |
Neural Network |
NOMA |
Non-Orthogonal Multiple Access |
NTN |
Non-Terrestrial Networks |
OAI |
OpenAirInterface |
OFDM |
Orthogonal Frequency Division Multiplexing |
O-RAN |
Open Radio Access Network |
OSI |
Open Systems Interconnection |
OSPF |
Open Shortest Path First |
PCA |
Principal Component Analysis |
PG |
Policy Gradient |
PGM |
Probabilistic Graph Model |
PNN |
Probabilistic Neural Network |
PSD |
Power Spectral Density |
PSO |
Particle Swarm Optimization |
QoE |
Quality of Experience |
QoS |
Quality of Service |
RC |
Reservoir Computing |
RAN |
Radio Access Network |
RBM |
Restricted Boltzmann Machine |
RF |
Random Forest |
RIC |
RAN Intelligent Controller |
RIS |
Reflective Intelligent Surface |
RL |
Reinforcement Learning |
RMS |
Root Mean Square |
RNN |
Recurrent Neural Network |
RSMA |
Rate-Splitting Multiple Access |
RSRP |
Reference Signal Received Power |
RSRQ |
Reference Signal Received Quality |
RU |
Radio Unit |
SA |
Simulated Annealing |
SAGIN |
Space-Air-Ground Integrated Networks |
SARSA |
State-Action-Reward-State-Action |
SDMA |
Space Division Multiple Access |
SDN |
Software Defined Network |
SDR |
Software Defined Radio |
SINR |
Signal to Interference and Noise Ratio |
SL |
Supervised Learning |
SNR |
Signal to Noise Ratio |
SoC |
System-on-Chip |
SOM |
Self-Organizing Map |
SpL |
Split Learning |
SVM |
Support Vector Machine |
TDD |
Time Division Duplexing |
TDMA |
Time Division Multiple Access |
TNTN |
integrated Terrestrial and Non-Terrestrial Network |
UE |
User Equipment |
UL |
Unsupervised Learning |
ULBC |
Ultra-reliable low Latency Broadband Communication |
uMBB |
ubiquitous Mobile BroadBand |
URLLC |
Ultra Reliable Low Latency Communication |
V2X |
Vehicle-to-Everything |
VR |
Virtual Reality |
VSAT |
Very Small Aperture Terminal |
WMMSE |
Weighted Minimum Mean Square Error |
@@ II-A Definition
Non-Terrestrial Network (NTN) refers to any network operating through the air or space-borne vehicle(s) for communication [19]. This definition implies that two distinct types of NTN platforms (space-borne and air-borne) can be utilized for NTN at different heights which is illustrated in Figure 2.
@@ II-A1 Space-borne platforms
Space-borne platforms, such as satellites, are deployed in space for communication [19]. They move around the Earth in specific orbits with varying angular velocities, relying on gravity to provide the necessary centripetal force to maintain their orbits. The orbital period of a satellite refers to the time required for the satellite to complete one full revolution around the Earth. Due to differences in orbital periods, some satellites may not be visible to ground observers all the time. To characterize this, another term is used to denote the duration of direct visibility for a satellite. This is known as the horizon time, which refers to the maximum duration during which the satellite is within the line of sight of a given ground station or receiver. Depending on their mobility with respect to the Earth, satellites can be classified into two broad categories: Geostationary (GEO/GSO) and Non-Geostationary (NGEO/NGSO) Earth Orbit satellites. We discuss these two types of satellites below and summarize their key features in Table II.
@@
Geostationary Earth Orbit (GEO or GSO) Satellites: These satellites have an orbital period of 24 hours which is the same as the time required for the Earth to complete a full rotation on its axis. As a result, these satellites appear stationary from the ground and are named Geostationary Earth Orbit (GEO or GSO) Satellites. These satellites orbit on the Earth’s equatorial plane at an altitude of about 35,786 km to maintain this orbital period. Due to this high altitude, it has an extremely large beam footprint (typically the diameter ranges from 200 to 1000 km) covering a pretty wide area. However, it also incurs an extremely long propagation delay (typically around 270 ms) [20] which makes it infeasible for low-latency communications. These satellites have been used in broadcasting services for a very long time, but are not very suitable for low-latency emerging applications.
@@
Non-Geostationary Earth Orbit (NGEO or NGSO) Satellites: As the name suggests, these satellites orbit around the Earth at a period lower than 24 hours, so they are not stationary with respect to a ground observer. As the orbital period is smaller, their angular velocity is also higher but the altitude is lower compared to GEO satellites. Depending on the heights, they can be divided into two categories: Low Earth Orbit (LEO) and Medium Earth Orbit (MEO) satellites. Typically they are deployed at a height ranging from 200 to 2000 km for LEO and 2000 to 25000 km for MEO satellites. The horizon time is much smaller for NGEO satellites due to smaller orbital periods, for example, the LEO satellites deployed at a height of around 500-600 km with an orbital period of 1.5-2 hours can have a horizon time of 5-10 minutes depending on channel conditions. Due to smaller heights, these satellites have a smaller beam footprint (diameter ranges from 5 to 500 km) with a much smaller propagation delay (typically around 20 ms for LEO satellites and 94 ms for MEO satellites)[20] compared to GEO satellites. With their proximity to Earth and lower cost of launch and maintenance, these satellites, especially the LEO satellites, have gained significant attention in recent years. Their reduced propagation delay and path loss make them an attractive choice for facilitating high-speed data transfer and real-time communication, so as to transform the future 6G connectivity.
@@
Orbital height (km) |
35786 |
2000-25000 |
200-2000 |
Typical diameter of beam footprint (km) |
1000 |
500 |
100 |
Propagation delay (ms) |
270 |
94 |
12-25 |
Orbital period (hours) |
24 |
12 |
1.5-2 |
Horizon time |
24 hours |
1-2 hours |
5-10 minutes |

### v4

@@ Abstract
I Introduction
@@
The Third Generation Partnership Project (3GPP) has already started the standardization towards the 5th Generation (5G)-Advanced in Release 17 and 18 to facilitate its worldwide deployment[5G-adv-1, 5G-adv-2]. 5G-Advanced provides much higher data rates, lower latency, increased capacity, and more efficient spectrum utilization than any of its predecessors. It supports a wide range of applications encompassing all 5G use cases such as Ultra Reliable Low Latency Communications (URLLC), massive Machine Type Communication (mMTC), and enhanced Mobile Broadband (eMBB) communication with different Key Performance Indicator (KPI) requirements [5G-adv]. Nevertheless, future applications such as Augmented Reality (AR), Virtual Reality (VR), Tactile Internet, Holographic Type Communication (HTC), remote health and surgery, etc. require extremely high throughput, low latency, high reliability, and ubiquity at the same time which cannot be met with current technological standards [6g-1]. Consequently, the next-generation global wireless standard, namely, 6th Generation (6G) has become the current research focus for the industry and research community [6g-ppf].
@@
6G is expected to provide an extremely high data rate (peak data rate up to 1 Tbps and user-experienced data rate up to 10 Gbps, around 100 times higher than 5G), very low latency (in the order of s), high reliability (around 100 times better than 5G) and extreme coverage to support the diverse set of future applications [6g-huawei, 6g-eric, 6g-samsung]. Due to the limited coverage area and geographical constraints, it is not possible to guarantee ubiquitous connectivity with existing terrestrial-only network infrastructures. Non-Terrestrial Networks (NTNs), networks involving space and aerial platforms, can provide us with multicast opportunities over very large areas as well as can serve users even in remote areas or during times of natural calamities [6g-1]. Furthermore, the launching and maintenance costs for satellites have significantly decreased as they are deployed at lower heights (typically around 600 km). These satellites can provide much higher throughput and lower latency compared to legacy satellites and potentially can support different use cases of 6G. So NTN is considered to become one of the major technological enablers of future 6G networks visioning connectivity anywhere and anytime [6g-1, 6g-2, 6g-3, 6g-4]. Tech giants such as SpaceX Starlink, Amazon Kuiper, and OneWeb have already begun to invest billions of dollars in this field, reflecting its massive potential for future growth [6g-sat].
@@
Although NTN presents numerous potential benefits for the development of future 6G networks, it also entails several challenges that need to be addressed, primarily due to the unique characteristics of its mobility and propagation environments [NTN_1, NTN_2]. Due to the long distances between the space-borne Base Stations (BS) and the ground User Equipment (UE), the propagation delay is usually higher in NTN environments. Additionally, high-speed air or space-borne platforms necessitate modifications to existing handover and paging protocols, as well as introduce a significant Doppler shift in carrier frequencies. The large path loss also increases the minimum power requirement for reliable transmission, initiating the need for novel beam and resource allocation strategies. Spectrum sharing in the same frequency band with existing terrestrial or other services also requires further study in order to avoid interference between terrestrial and non-terrestrial users. Even though currently there are some stand-alone satellite network deployments, the ultimate goal is the convergence of terrestrial and non-terrestrial environments for extreme network performance in 6G. [6g-1]. This potential integrated environment requires efficient computing, routing, and slicing algorithms for meeting the expected KPI requirements of 6G.
@@
Artificial intelligence (AI) is currently having a profound and revolutionary impact on a multitude of industries, including but not limited to healthcare, military, transportation, and e-Commerce [ai]. AI encompasses a wide array of smart machines, while Machine Learning (ML) is a popular subset of AI that allows machines to learn from large amounts of data and make decisions without the need for explicit programming [ai-ml]. Deep Learning (DL) is a special subset of ML that studies Artificial Neural Networks (ANNs) which contain more than one hidden layer, often implemented to simulate the human brain [ai-dl]. DL is currently being leveraged in various applications, such as computer vision, speech recognition, and bioinformatics, outperforming human-level performance in these particular domains. The cellular domain is still in its infancy in terms of AI integration [ai-wireless] compared to other fields due to the complex and dynamic nature of wireless networks. As an integral part of 6G networks, challenges associated with NTN deployment provides an enticing field for AI applications. However, while deploying algorithms in a real environment, practical implementation difficulties may arise to provide reliable vertical connectivity between the ground and space networks. To reach optimal performance, theoretical advancements in communication system design must be complemented by appropriate AI solutions for NTN integration into 6G.
@@
| 3GPP | Third Generation Partnership Project |
| 5G | 5th Generation |
| 6G | 6th Generation |
| AC | Actor-Critic |
| ACO | Ant-Colony based Optimization |
| AI | Artificial Intelligence |
| ANN | Artificial Neural Networks |
| AoA | Angle of Arrival |
| AoD | Angle of Departure |
| AR | Augmented Reality |
| ARIMA | Auto-Regressive Moving Average |
| ARMA | Auto-Regressive Integrated Moving Average |
| ATM | Asynchronous Transfer Mode |
| BS | Base Station |
| CCI | Co-Channel Interference |
| CDMA | Code Division Multiple Access |
| CG | Coordinate Graph |
| CNN | Convolutional Neural Network |
| CSD | Cyclo-Stationary Detection |
| CSI | Channel State Information |
| CU | Central Unit |
| DBN | Deep Belief Network |
| DcL | Decentralized Learning |
| DDPG | Deep Deterministic Policy Gradient |
| DDQN | Double Deep Q-Learning Network |
| DL | Deep Learning |
| DN | Deconvolutional Network |
| DoS | Denial-of-Service |
| DP | Dynamic Programming |
| DPG | Deterministic Policy Gradient |
| DQN | Deep Q-Learning Network |
| DRL | Deep Reinforcement Learning |
| DRN | Deep Residual Network |
| DRQN | Deep Recurrent Q-Learning Network |
| DU | Distributed Unit |
| ED | Energy Detection |
| eMBB | enhanced Mobile Broadband |
| ELM | Extreme Learning Machine |
| ESA | European Space Agency |
| ESN | Echo-State Network |
| EVD | Eigen Value-based Detection |
| FCNN | Fully Connected Neural Network |
| FDMA | Frequency Division Multiple Access |
| FL | Federated Learning |
| FlexRIC | Flexible RIC |
| GA | Genetic Algorithm |
| GDM | Generative Diffusion Model |
| GMM | Gaussian Mixture Model |
| GAN | Generative Adversarial Network |
| GNN | Graph Neural Network |
| GEO/GSO | Geostationary Earth Orbit |
| GNSS | Global Navigation Satellite System |
| GPS | Global Positioning System |
| GRU | Gated Recurrent Unit |
| HAPS | High Altitude Platform System |
| HARQ | Hybrid Automatic Repeat Request |
| HIBS | High-altitude International Mobile Base Station |
| HTC | Holographic Type Communication |
| IC | Integrated Circuit |
| ICI | Inter-Carrier Interference |
| IoT | Internet of Things |
| IP | Internet Protocol |
| ISL | Inter-Satellite Link |
| ITU | International Telecommunication Union |
| KKT | KarushKuhnTucker |
| KNN | k-Nearest Neighbor |
| KPI | Key Performance Indicator |
| LEO | Low Earth Orbit |
| LoS | Line of Sight |
| LSM | Liquid State Machine |
| LSTM | Long-Term Short Memory |
@@
| MADRL | Multi-Agent Deep Reinforcement Learning |
| MAP | Maximum A Posteriori |
| MARL | Multi-Agent Reinforcement Learning |
| MC | Monte Carlo method |
| MDP | Markov Decision Process |
| MEMS | Micro-Electro-Mechanical Systems |
| MEO | Medium Earth Orbit |
| MiM | Man-in-the-Middle |
| ML | Machine Learning |
| MLP | Multi-Layer Perceptron |
| MMSE | Minimum Mean Square Error |
| mMTC | massive Machine Type Communication |
| MNL | Minimum Network Load |
| MSQ | Maximum Signal Quality |
| MST | Maximum Service Time |
| mULC | massive Ultra-reliable low-Latency Communication |
| NGEO/NGSO | Non-Geostationary Earth Orbit |
| NN | Neural Network |
| NOMA | Non-Orthogonal Multiple Access |
| NTN | Non-Terrestrial Networks |
| OAI | OpenAirInterface |
| OFDM | Orthogonal Frequency Division Multiplexing |
| O-RAN | Open Radio Access Network |
| OSI | Open Systems Interconnection |
| OSPF | Open Shortest Path First |
| PCA | Principal Component Analysis |
| PG | Policy Gradient |
| PGM | Probabilistic Graph Model |
| PNN | Probabilistic Neural Network |
| PSD | Power Spectral Density |
| PSO | Particle Swarm Optimization |
| QoE | Quality of Experience |
| QoS | Quality of Service |
| RC | Reservoir Computing |
| RAN | Radio Access Network |
| RBM | Restricted Boltzmann Machine |
| RF | Random Forest |
| RIC | RAN Intelligent Controller |
| RIS | Reflective Intelligent Surface |
| RL | Reinforcement Learning |
| RMS | Root Mean Square |
| RNN | Recurrent Neural Network |
| RSMA | Rate-Splitting Multiple Access |
| RSRP | Reference Signal Received Power |
| RSRQ | Reference Signal Received Quality |
| RU | Radio Unit |
| SA | Simulated Annealing |
| SAGIN | Space-Air-Ground Integrated Networks |
| SARSA | State-Action-Reward-State-Action |
| SDMA | Space Division Multiple Access |
| SDN | Software Defined Network |
| SDR | Software Defined Radio |
| SINR | Signal to Interference and Noise Ratio |
| SL | Supervised Learning |
| SNR | Signal to Noise Ratio |
| SoC | System-on-Chip |
| SOM | Self-Organizing Map |
| SpL | Split Learning |
| SVM | Support Vector Machine |
| TDD | Time Division Duplexing |
| TDMA | Time Division Multiple Access |
| TNTN | integrated Terrestrial and Non-Terrestrial Network |
| UE | User Equipment |
| UL | Unsupervised Learning |
| ULBC | Ultra-reliable low Latency Broadband Communication |
| uMBB | ubiquitous Mobile BroadBand |
| URLLC | Ultra Reliable Low Latency Communication |
| V2X | Vehicle-to-Everything |
| VR | Virtual Reality |
| VSAT | Very Small Aperture Terminal |
| WMMSE | Weighted Minimum Mean Square Error |
@@ II-A Definition
Non-Terrestrial Network (NTN) refers to any network operating through the air or space-borne vehicle(s) for communication [5G-space]. This definition implies that two distinct types of NTN platforms (space-borne and air-borne) can be utilized for NTN at different heights which is illustrated in Figure 2.
@@ II-A1 Space-borne platforms
Space-borne platforms, such as satellites, are deployed in space for communication [5G-space]. They move around the Earth in specific orbits with varying angular velocities, relying on gravity to provide the necessary centripetal force to maintain their orbits. The orbital period of a satellite refers to the time required for the satellite to complete one full revolution around the Earth. Due to differences in orbital periods, some satellites may not be visible to ground observers all the time. To characterize this, another term is used to denote the duration of direct visibility for a satellite. This is known as the horizon time, which refers to the maximum duration during which the satellite is within the line of sight of a given ground station or receiver. Depending on their mobility with respect to the Earth, satellites can be classified into two broad categories: Geostationary (GEO/GSO) and Non-Geostationary (NGEO/NGSO) Earth Orbit satellites. We discuss these two types of satellites below and summarize their key features in Table II.
@@
Geostationary Earth Orbit (GEO or GSO) Satellites: These satellites have an orbital period of 24 hours which is the same as the time required for the Earth to complete a full rotation on its axis. As a result, these satellites appear stationary from the ground and are named Geostationary Earth Orbit (GEO or GSO) Satellites. These satellites orbit on the Earth’s equatorial plane at an altitude of about 35,786 km to maintain this orbital period. Due to this high altitude, it has an extremely large beam footprint (typically the diameter ranges from 200 to 1000 km) covering a pretty wide area. However, it also incurs an extremely long propagation delay (typically around 270 ms) [3gpp1] which makes it infeasible for low-latency communications. These satellites have been used in broadcasting services for a very long time, but are not very suitable for low-latency emerging applications.
@@
Non-Geostationary Earth Orbit (NGEO or NGSO) Satellites: As the name suggests, these satellites orbit around the Earth at a period lower than 24 hours, so they are not stationary with respect to a ground observer. As the orbital period is smaller, their angular velocity is also higher but the altitude is lower compared to GEO satellites. Depending on the heights, they can be divided into two categories: Low Earth Orbit (LEO) and Medium Earth Orbit (MEO) satellites. Typically they are deployed at a height ranging from 200 to 2000 km for LEO and 2000 to 25000 km for MEO satellites. The horizon time is much smaller for NGEO satellites due to smaller orbital periods, for example, the LEO satellites deployed at a height of around 500-600 km with an orbital period of 1.5-2 hours can have a horizon time of 5-10 minutes depending on channel conditions. Due to smaller heights, these satellites have a smaller beam footprint (diameter ranges from 5 to 500 km) with a much smaller propagation delay (typically around 20 ms for LEO satellites and 94 ms for MEO satellites)[3gpp1] compared to GEO satellites. With their proximity to Earth and lower cost of launch and maintenance, these satellites, especially the LEO satellites, have gained significant attention in recent years. Their reduced propagation delay and path loss make them an attractive choice for facilitating high-speed data transfer and real-time communication, so as to transform the future 6G connectivity.
@@
| Orbital height (km) | 35786 | 2000-25000 | 200-2000 |
| Typical diameter of beam footprint (km) | 1000 | 500 | 100 |
| Propagation delay (ms) | 270 | 94 | 12-25 |
| Orbital period (hours) | 24 | 12 | 1.5-2 |
| Horizon time | 24 hours | 1-2 hours | 5-10 minutes |